OpenStudy (anonymous):

Simplify the expression (2x − 9)(x + 6)

OpenStudy (anonymous):

do you know how to foil?

OpenStudy (michele_laino):

please, you have to apply the distributive property of multiplication over algebraic sum!

OpenStudy (anonymous):

I think the foil method would be best in this case. Foil stands for First Outer Inner Last. So you would take the first numbers in the equation and multiply. \[2x \times x= 2x ^{2}\] the outer and inner steps go together. You multiply the outer numbers first: \[2x \times 6 = 12x\] Then, the inner numbers: \[-9 \times x =-9x\] Then you would add both of the products: \[12x - 9x = 3x\] Now, you will multiply the last numbers. \[-9 \times 6 =-54\] So your answer will be: \[2x ^{2}+3x -54\] F O I L Hope this helps! :)

OpenStudy (michele_laino):

that's right!